# Wave Function Collapse, Lorentz Invariance, and the Third Postulate of Relativity

### v3

@@ Abstract
The changes that quantum states undergo during measurement are both probabilistic and nonlocal. These two characteristics complement one another to insure compatibility with relativity and maintain conservation laws. Nonlocal entanglement relations provide a means to enforce conservation laws in a probabilistic theory, while the probabilistic nature of nonlocal effects prevents the superluminal transmission of information. In order to explain these measurement-induced changes in terms of fundamental physical processes it is necessary to take these two key characteristics into account. One way to do this is to modify the Schrödinger equation by adding stochastic, nonlinear terms. A number of such proposals have been made over the past few decades. A recently proposed equation based on the assumption that wave function collapse is induced by a sequence of correlating interactions of the kind that constitute measurements has been shown to maintain strict adherence to conservation laws in individual instances, and has also eliminated the need to introduce any new, ad hoc physical constants.
In this work it is shown that the stochastic modification to the Schrödinger equation is Lorentz invariant. It is further argued that the additional spacetime structure that it requires provides a way to implement the assumption that spacelike-separated operators (and measurements) commute, and that this assumption of local commutativity should be regarded as a third postulate of relativity.
@@ I Introduction
In order to provide a fully coherent logical framework for contemporary theory it is crucial that any proposed modifications to the fundamental equations be Lorentz invariant. The goals of this work are to show that the proposal of [16] meets this criterion, and to examine its implications for our understanding of relativity. The next section describes the basic assumptions employed by nonlinear, norm-preserving stochastic collapse equations and illustrates how they work. Section III reviews the proposed modification, showing how it eliminates the need to introduce new ad hoc physical constants and insures that conservation laws are respected in individual experiments. It then goes on to demonstrate that it meets a critical test of Lorentz invariance in a very natural way. Based on the demonstration of Lorentz invariance Section IV examines the implications for our understanding of relativity and spacetime ontology.
@@ II.2 Nonlinear, Norm-Preserving Stochastic Collapse Equations
Measurements of quantum systems can have a very large number of possible outcomes, but at the most basic level they come down to either a detection or a failure to detect. The essentially binary character of measurement processes means that at each stage the Hilbert space (of any number of dimensions) can be decomposed into two orthogonal subspaces, and the state vector of the total system can be represented as the sum of two components, one in each subspace. Although the particular decomposition can vary during the process (thus, allowing any number of possible outcomes) the fundamentally binary nature makes it possible to model measurement processes as a random walk between two alternatives. [^2] The claim that measurement processes are essentially binary might seem implausible when one considers many common-place measurements that we make such as length, height, weight, etc. However, that is because we typically describe these kinds of measurements by reference to extremely complex human actions. Imagine eliminating any reference to humans and try to design a robot to carry out the measurement. Eventually one must get back to describing the fundamental physical processes involved. This is essentially what stochastic collapse equations do. [^3] Wave function collapse can occur either inside or outside a laboratory, but for ease of explanation the discussion here will focus primarily on typical measurement situations.
[…]
The binary character of measurement processes noted above allows us to illustrate the essential operation of stochastic collapse equations using a simple two state system. Consider a system with a wave function $\psi$ and Hamiltonian, $\mathbf{H}$. The system evolves under the action of the Hamiltonian and of a stochastic collapse operator that is constructed from a self-adjoint operator, $\mathbf{O}$, with two eigenstates, $|x\rangle$ and $|y\rangle$, associated with eigenvalues $\mathbf{a}$ and $\mathbf{b}$. The eigenstates of the operator, $\mathbf{O}$, define the collapse basis. The nonlinear stochastic operator is defined as: $\,\mathcal{O}\;\equiv\;k[\mathbf{O}-\langle\,\psi|\mathbf{O}|\psi\,\rangle]\,$, where $\langle\,\psi|\mathbf{O}|\psi\,\rangle$ is the expectation value of the operator, $\mathbf{{O}}$, in the state, $\psi$, and $k$ is a constant that helps to determine the strength and scale of the collapse effects and also insures that $\mathcal{O}$ is dimensionless. For example, if $\mathcal{O}$ is based on the position operator $k$ could determine the range of the collapse effect.
@@
[…]
@@
In this form and with the simplifying assumptions described above it is possible to trace the evolution of the wave function through Hilbert space under the influence of the stochastic operator. Since $\alpha$ and $\beta$ are assumed to be real and positive the evolution can be modeled as a random walk along the arc joining $x$ and $y$ axes (corresponding to the eigenstates $|x\rangle$ and $|y\rangle$). [^4] In this simplified picture with $\alpha$ and $\beta$ real and positive the process can be pictured on a standard Cartesian graph with x and y axes. It is very straightforward to transfer the analysis to the Bloch sphere afterward. The state, $\psi\,=\,\alpha|x\rangle\,+\,\beta|y\rangle$, lies on the arc; the orthogonal state, $\beta|x\rangle\,-\,\alpha|y\rangle$, is tangent to the arc and it drives the state to one or the other of the eigenstates in infinitesimal steps. The magnitude and direction of the steps are determined by the coefficient of the tangent state vector, $\alpha\,\beta\;k(\mathbf{a}-\mathbf{b})\sqrt{\gamma}\,d\xi(t).$
The dependence on $\alpha$ and $\beta$ means that the step size varies as the wave function, $\psi$, traverses the arc under the influence of the stochastic operator. Therefore, the calculation of the distance to the end points, $|x\rangle$ and $|y\rangle$, must take this variation into account. This can be done by introducing a parameter, $\theta$, with $\alpha\,=\,\cos\,\theta$ and $\beta\,=\,\sin\,\theta$, and integrating the coefficient of the tangent state vector along the arc:
@@
[…]
@@ III Interaction-Induced Wave Function Collapse
The rate can be defined so that it integrates to a value of approximately, $1$, over the course of the interaction, and goes to zero when the interacting systems either settle into a stationary state or separate to a distance at which the interaction effectively ends. This insures that the stochastic collapse equation reduces to the ordinary Schrödinger equation in these situations. This can be done by taking the ratio of the rate of change of potential energy to the maximum magnitude of potential energy that occurs during the interaction:
@@
$\begin{array}[]{ll}\Big{|}\Big{|}\;(i\hbar)\,\mathbf{\int}\,\mathbf{{V}_{jk}}\,\Big{(}\,\frac{\psi^{*}\,\mathbf{\nabla_{j}}^{2}\psi\,-\,\psi\mathbf{\nabla_{j}}^{2}\psi^{*}}{2m_{j}}\,+\,\frac{\psi^{*}\,\mathbf{\nabla_{k}}^{2}\psi\,-\,\psi\mathbf{\nabla_{k}}^{2}\psi^{*}}{2m_{k}}\,\Big{)}\;\Big{|}\Big{|}.\end{array}$ (6)
@@
To establish that the expression for the denominator in $\gamma_{jk}$, $||\langle\,\psi|\mathbf{{V}_{jk}}|\,\psi\rangle||_{max}$, is well defined note that the collapse equation can be integrated both backward and forward in time. Because the Hamiltonian is independent of time it is clear that the associated Schrödinger equation is integrable. The fact that the stochastic operator is unbiased implies that averaging over all of the unravellings of the collapse equation will reproduce the value obtained by the Schrödinger equation integral. Observe also that the stochastic changes in the denominator are matched by those in the numerator; so their overall effect on $\gamma_{jk}$ is to leave it essentially unchanged. It is, of course, not necessary to actually carry out the integration since $\gamma_{jk}$ is designed to integrate to a value of order, $1$, over the course of the interaction. With this understanding the variable rate parameter is defined as follows:
$\gamma_{jk}\,\equiv\,\frac{||(i\hbar)\,\mathbf{\int}\,\mathbf{{V}_{jk}}\Big{(}\frac{\psi^{*}\mathbf{\nabla_{j}}^{2}\psi-\psi\mathbf{\nabla_{j}}^{2}\psi^{*}}{2m_{j}}+\frac{\psi^{*}\mathbf{\nabla_{k}}^{2}\psi-\psi\mathbf{\nabla_{k}}^{2}\psi^{*}}{2m_{k}}\,\Big{)}||}{||\langle\,\psi|\mathbf{{V}_{jk}}|\,\psi\rangle||_{max}}.$ (7)
An important point to note about 7 is that both numerator and denominator pick out only the interacting components of the the wave function. This insures that $\gamma_{jk}$ is dependent only on the portion of the wave function that is involved in the interaction. This insures that the rate parameter is independent of the amplitudes of the interacting and noninteracting components. This feature is necessary to insure compliance with the Born probability rule as illustrated in the discussion in Section II.
@@
In discussing the status of conservation laws with reference to stochastic collapse equations there are two issues that must be addressed. First, some of the collapse proposals imply small violations of conservation laws even with regard to statistical averages. For example, consider the most widely known collapse proposal, the continuous spontaneous localization (CSL) model[36, 37]. It implies small violations of energy conservation as described in [38]. One of the main reasons for these violations is that this proposal uses a stochastic field, $\xi(x,t)$, which is a function of both position and time, as opposed to a single, global stochastic process, $\xi(t)$, which is a function only of time. The stochastic field induces spatial variations in the wave function - in effect, pumping energy into it. Other proposals, such as one of the variants discussed in [1] use a single, global stochastic process, and bases the stochastic operator on the Hamiltonian. In this way it maintains energy conservation on average, but still violates it in individual instances.
Because the proposed equation in [16] uses a single global stochastic process it avoids the kinds of statistical violations of conservation laws just described. The reason that it is able to maintain strict conservation in individual measurement situations is that it is based on the assumption that the amplitude shifts between branches that bring about collapse are generated by conservative interactions. So these interactions are responsible both for the splitting of the wave function into branches and for the redistribution of amplitude among the branches. As stated above subsequent to the splitting under ordinary Schrödinger evolution conserved quantities are identical in each branch (up to normalization of the branch). Thus, the amplitude shifts between branches do not alter any of these quantities, and the eventual selection (or elimination) of a particular branch leaves the surviving wave function with the same values that it had prior to the chain of interactions (including preparation) that led to the collapse.
It is shown in [16] that the proposed equation conserves momentum and angular momentum exactly. Because the proposal is formulated in a nonrelativistic framework it is only able to conserve energy within the accuracy allowed by the limited forms of energy describable in nonrelativistic theory.
@@
This review of the proposal in [16] is intended as background for the demonstration that the action of the stochastic operator in 10 is Lorentz invariant. The general question of Lorentz invariance hangs over any proposal that deals with the measurement problem. But there is a particular concern that arises in connection with collapse equations of the type discussed here because they are formulated in a preferred reference frame. There is essentially zero probability that any reference frame in which one chooses to analyze a collapse process will coincide with the preferred frame. The critical question is whether the magnitude of the collapse effects is independent of the frame in which they are considered.
The sort of discrepancies in predictions of correlations mentioned two paragraphs back occur in any nonlinear collapse equation. (Other experimental deviations from conventional theory may also occur.) If the predicted magnitude of the collapse effects differs from one frame to the next there will be, in principle, observable differences from the preferred frame and the proposal will be in explicit conflict with relativity. So the task here is to show that the magnitude of the collapse effects predicted by equation 10 is independent of the reference frame in which they are viewed.
According to equation 10 during the brief time in which an interaction generates a correlation between two systems it also transfers amplitude between the interacting and noninteracting branches of the wave function. The magnitude of the transfer is:
${\mathcal{V}}_{jk}\sqrt{\gamma_{jk}}d\xi(t)\,\psi\;\equiv\;\frac{\mathbf{{V}_{jk}}-\langle\,\psi|\mathbf{{V}_{jk}}|\psi\,\rangle}{(m_{j}+m_{k})c^{2}}\,\sqrt{\gamma_{jk}}d\xi(t)\,\psi.$ (11)
We want to show that this magnitude is the same whether it is calculated in the rest frame of the laboratory in which the interaction takes place or in the preferred frame in which the collapse equation is formulated.
Consider, first, the term, $\sqrt{\gamma_{jk}}d\xi(t)$, that governs the rate at which the transfer takes place. As argued above, this term integrates to a value of order, $1$, over the course of the interaction. As viewed from the preferred frame the duration of the interaction is increased by the relativistic time dilation factor, and the rate parameter, $\gamma_{jk}$, is reduced by the same proportion. But, from the way in which the rate parameter is constructed it is clear that it will still integrate to the same value over the course of the interaction. So we need only show that the expression, $\frac{\mathbf{{V}_{jk}}-\langle\,\psi|\mathbf{{V}_{jk}}|\psi\,\rangle}{(m_{j}+m_{k})c^{2}}$ is invariant.
To demonstrate this it is necessary to show that the relativistic transformation of the numerator is the same as that of the denominator. The numerator is based on the interaction potential energy. The elementary interactions that occur in nonrelativistic quantum theory are electromagnetic. Therefore, the potential energy is an electromagnetic potential. From a relativistic point of view this potential is the timelike component of an electromagnetic four-potential, and transforms accordingly. The denominator consists of the sum of the relativistic energies of the interacting systems. The relativistic energy is the timelike component of the energy-momentum four vector. So the denominator transforms in the same way as the numerator, and, hence, the expression for the collapse term is Lorentz invariant.

### v4

@@ Abstract
The changes that quantum states undergo during measurement are both probabilistic and nonlocal. These two characteristics complement one another to insure compatibility with relativity and maintain conservation laws. The probabilistic nature of nonlocal effects prevents the superluminal transmission of information, while nonlocal entanglement relations provide a means to enforce conservation laws in a probabilistic theory. In order to explain measurement-induced changes in terms of fundamental physical processes it is important to take these two key characteristics into account. One way to do this is to modify the Schrödinger equation by adding stochastic, nonlinear terms. A number of such proposals have been made over the past few decades. A recently proposed equation based on the assumption that wave function collapse is induced by a sequence of correlating interactions of the kind that constitute measurements has been shown to maintain strict adherence to conservation laws in individual instances, and has also eliminated the need to introduce any new, ad hoc physical constants.
In this work it is shone that the proposed stochastic modification to the Schrödinger equation is also Lorentz invariant, even though it is formulated in a preferred reference frame. It is further argued that the additional spacetime structure that the proposed modification requires provides a way to implement the assumption that spacelike-separated operators (and measurements) commute, and that this assumption of local commutativity should be regarded as a third postulate of relativity.
@@ I Introduction
In order to provide a fully coherent logical framework for contemporary theory it is crucial that any proposed modifications to the fundamental equations be Lorentz invariant. The goals of this work are to show that the proposal of [16] meets this criterion, and to examine its implications for our understanding of relativity. The next section describes the basic assumptions employed by nonlinear, norm-preserving stochastic collapse equations and illustrates how they work. Section III reviews the modification proposed in [16], showing how it eliminates the need to introduce new ad hoc physical constants and insures that conservation laws are respected in individual experiments. Section IV demonstrates that the stochastic modification described in the previous section is Lorentz invariant, despite being formulated in a nonrelativistic framework. Based on the demonstration of Lorentz invariance Section V examines the implications for our understanding of relativity and spacetime ontology.
@@ II.2 Nonlinear, Norm-Preserving Stochastic Collapse Equations
Measurements of quantum systems can have a very large number of possible outcomes, but at the most basic level they come down to either a detection or a failure to detect. The essentially binary character of measurement processes means that at each stage the Hilbert space (of any number of dimensions) can be decomposed into two orthogonal subspaces, and the state vector of the total system can be represented as the sum of two components, one in each subspace. Although the particular decomposition can vary during the process (thus, allowing any number of possible outcomes) the fundamentally binary nature makes it possible to model measurement processes as a random walk between two alternatives. [^2] When there are more than two possible outcomes the alternatives consist of two disjoint sets of outcomes, one of which will increase in amplitude and the other decrease. This is essentially what stochastic collapse equations do. [^3] Wave function collapse can occur either inside or outside a laboratory, but for ease of explanation the discussion here will focus primarily on typical measurement situations.
To see how these equations work we can begin with a simple theorem about the relative probabilities of a random walk ending at either of the two end points. If the walk continues for enough steps it will eventually finish at one or the other of the end points. Label the end points as 0 and $1$, and suppose that the walk begins at a point, $p$, between them. It will be shown that the probability of ending at $1$ is $p$ and that the probability of ending at 0 is 1p1-p. The step size is labeled $\delta$. It is allowed to vary anywhere between 0 and the distance to the nearest end point: $0\,\leq\,\delta\,\leq\,p$, $0\,\leq\,\delta\,\leq\,1-p$. Label the probability of reaching $1$ as $Pr(p)$. Because the walk is assumed to be unbiased we get: $Pr(p)\,=\,\frac{1}{2}Pr(p-\delta)+\frac{1}{2}Pr(p+\delta)$. Because this relationship holds for all values of $p$ and $\delta$ it is linear. Therefore, the probability of reaching $1$ is $p$, and the probability of reaching 0 is 1p1-p. What follows is an attempt to present an intuitive explanation of how nonlinear, norm-preserving stochastic collapse equations map onto this simple, binary picture.
The binary character of measurement processes noted above allows us to illustrate the essential operation of stochastic collapse equations using a simple two state system. Consider a system with a wave function, $\psi$, and Hamiltonian, $\mathbf{H}$. The system evolves under the action of the Hamiltonian and of a stochastic collapse operator that is constructed from a self-adjoint operator, $\mathbf{O}$, with two eigenstates, $|x\rangle$ and $|y\rangle$, associated with eigenvalues $\mathbf{a}$ and $\mathbf{b}$. The eigenstates of the operator, $\mathbf{O}$, define the collapse basis. The nonlinear stochastic operator is defined as: $\mathcal{O}\;\equiv\;k[\mathbf{O}-\langle\,\psi|\mathbf{O}|\psi\,\rangle]\,$, where $\langle\,\psi|\mathbf{O}|\psi\,\rangle$ is the expectation value of the operator, $\mathbf{{O}}$, in the state, $\psi$, and $k$ is a constant that helps to determine the strength and scale of the collapse effects and also insures that $\mathcal{O}$ is dimensionless. For example, if $\mathcal{O}$ is based on the position operator $k$ could determine the range of the collapse effect.
@@
$\psi\,=\,\alpha|x\rangle+\beta|y\rangle$, with $|\alpha|^{2}\,+\,|\beta|^{2}\,=\,1$. To simplify the example these amplitudes can be taken as real and positive with no loss of generality. The action of the self-adjoint operator, $\mathbf{O}$, on the wave function is:
@@
In this form and with the simplifying assumptions described above it is possible to trace the evolution of the wave function through Hilbert space under the influence of the stochastic operator. Since $\alpha$ and $\beta$ are assumed to be real and positive the evolution can be modeled as a random walk along the arc joining $x$ and $y$ axes (corresponding to the eigenstates $|x\rangle$ and $|y\rangle$). [^4] In this simplified picture with $\alpha$ and $\beta$ real and positive the process can be pictured on a standard Cartesian graph with x and y axes. It is very straightforward to transfer the analysis to the Bloch sphere afterward. The state, $\psi\,=\,\alpha|x\rangle\,+\,\beta|y\rangle$, lies on the arc; the orthogonal state, $\beta|x\rangle\,-\,\alpha|y\rangle$, is tangent to the arc and it drives the state to one or the other of the eigenstates in infinitesimal steps. The magnitude and direction of the steps are determined by the coefficient of the tangent state vector, $\alpha\,\beta\;k(\mathbf{a}-\mathbf{b})\sqrt{\gamma}\,d\xi(t).$  [^5] The role of the third term in equation 1 involving $\mathcal{O}^{2}$ is simply to keep the resulting state on the arc by maintaining the correct normalization.
The dependence on $\alpha$ and $\beta$ means that the step size varies as the wave function, $\psi$, traverses the arc under the influence of the stochastic operator. Since the effective distance to the end points, $|x\rangle$ and $|y\rangle$, is measured in terms of this step size the distance calculation must take this variation into account. This can be done by introducing a parameter, $\theta$, with $\alpha\,=\,\cos\,\theta$ and $\beta\,=\,\sin\,\theta$, and integrating the coefficient of the tangent state vector along the arc:
@@
where $A$ is some term independent of $\alpha,\beta$ and $\theta$. If we associate the two eigenstates with the end points of a random walk as described earlier, with $|x\rangle$ corresponding to 0 and $|y\rangle$ corresponding to $1$, then the position of $\psi$ along the arc can be parameterized as $\beta^{2}$. As shown earlier, this is the probability that the random walk ends at $|y\rangle$, and $\alpha^{2}\,=\,1-\beta^{2}$ is the probability that it ends at $|x\rangle$. So the Born rule follows from the basic structure of the collapse equation in a straightforward manner. It is also worth noting that the random walk terminates because the term, $\alpha\beta\,=\cos\,\theta\,\sin\,\theta$ goes to 0 as $\psi$ approaches one of the end points. [^6] The fact that $\alpha\,\beta$ approaches 0 at the end points also creates a problem (the “tails problem”) in that the walk does not end in a finite number of steps. This problem will not be dealt with in detail here, but I will offer a speculative solution later.
@@ III Interaction-Induced Wave Function Collapse
The rate can be defined so that it integrates to a value of order, $1$, over the course of the interaction, and goes to zero when the interacting systems either settle into a stationary state or separate to a distance at which the interaction effectively ends. This insures that the stochastic collapse equation reduces to the ordinary Schrödinger equation in these situations. This can be done by taking the ratio of the rate of change of potential energy to the maximum magnitude of potential energy that occurs during the interaction:
@@
[…]
@@
To establish that the expression for the denominator in $\gamma_{jk}$, $\langle\,\psi|\mathbf{{V}_{jk}}|\,\psi\rangle||_{max}$, is well defined note that the collapse equation can be integrated both backward and forward in time. Because the Hamiltonian is independent of time it is clear that the associated Schrödinger equation is integrable. The fact that the stochastic operator is unbiased implies that averaging over all of the unravellings of the collapse equation will reproduce the value obtained by the Schrödinger equation integral. Observe also that the stochastic changes in the denominator are matched by those in the numerator; so their overall effect on $\gamma_{jk}$ is to leave it essentially unchanged. It is, of course, not necessary to actually carry out the integration since $\gamma_{jk}$ is designed to integrate to a value of order, $1$, over the course of the interaction. With this understanding the variable rate parameter is defined as follows:
[…]
An important point to note about 7 is that both numerator and denominator pick out only the interacting components of the the wave function. This means that $\gamma_{jk}$ depends only on the portion of the wave function that is involved in the interaction, and thus the rate parameter is independent of the amplitudes of the interacting and noninteracting components. This feature is necessary to insure compliance with the Born probability rule as illustrated in the discussion in Section II.
@@
The collapse process involves the shift of amplitude among distinct branches of the wave function. Because the total amount of a conserved quantity up to normalization is the same in each branch these shifts do not, by themselves, violate any conservation law.
Violations could occur if they are induced by the process that generates the shifts. There are a couple of ways in which such violations can be induced for the general type of collapse equations considered here. One of these is illustrated by the most widely known collapse proposal, the continuous spontaneous localization (CSL) model[9, 13], which implies small violations of energy conservation as described in [36].
The CSL proposal introduces two new constants and a stochastic field, $\xi(x,t)$, which is a function of both position and time. The new constants together with the stochastic variations in both space and time act as a source of external energy that can be fed into other physical systems. In contrast, the proposed equation in [16] does not introduce any new constants and it uses a single global stochastic process, $\xi(t)$, which is a function only of time. Therefore, this problem does not arise.
Another potential problem concerns the use of a stochastic process based on white noise. Since white noise includes arbitrarily high frequencies it will introduce energy violations if it is not modulated. To avoid this a number of proposals have employed colored noise, in which the high frequencies are strongly attenuated. In the current proposal the stochastic differential, $d\xi(t)$, is multiplied by the variable interaction rate, $\sqrt{\gamma_{jk}}$. The term, $\sqrt{\gamma_{jk}}d\xi(t)$, is significantly different from zero for only a very brief time, and it heavily suppresses frequencies that are much different from $\frac{\mathbf{{V}_{jk-max}}}{\hbar}$ (the inverse of equation 8). This prevents spurious high frequency noise from injecting additional energy into the system.
Explicit demonstrations of how the proposed equation maintains conservation laws in individual measurement situations are given in in [16]. Momentum and angular momentum are conserved exactly. Because the proposal is formulated in a nonrelativistic framework it is only able to conserve energy within the accuracy allowed by the limited forms of energy describable in nonrelativistic theory. Specifically, it is shown that the proposed modification exactly conserves energy in those situations in which the Schrödinger equation correctly predicts energy conservation. In cases where the Schrödinger equation fails to track relativistic adjustments to the nonrelativistic energy formulas the differences from those formulas that the stochastic modification predicts match the relativistic corrections both qualitatively and quantitatively.
@@
This review of the proposal in [16] is intended as background for the demonstration that the action of the stochastic operator in 10 is Lorentz invariant. The argument is presented in the next section.
IV Lorentz Invariance
To establish Lorentz invariance it must be shown that the amplitude shift induced by the interaction of two physical systems is independent of the reference frame from which that interaction is viewed. As described in the previous section the magnitude of the amplitude shift is determined by the ratio of the interaction energy to the total relativistic energy of the interacting systems. The stochastic operator associated with the shift is active only during the very brief period during which conserved quantities are being exchanged. The rate parameter is determined by the rate of change of interaction energy, and when multiplied by the stochastic differential in the appropriate way it integrates to a value of order one over the course of the interaction. It will be shown here that both the ratio of the interaction energy to the total relativistic energy of the two systems and the integration of the interaction rate over the course of the correlating phase of the interaction are independent of the reference frame.
The total relativistic energy of interacting systems, $j$ and $k$, consists of the rest energy and kinetic energy of the two systems, along with the interaction (aka “potential”) energy:
$\Big(\frac{m_{j}}{\sqrt{1-\beta_{j}^{2}}}+\frac{m_{k}}{\sqrt{1-\beta_{k}^{2}}}\Big)c^{2}\,+\mathbf{V_{jk}}.$ (11)
In this expression $m_{j}$ and $m_{k}$ represent the rest mass of the two systems. [^7] In the earlier discussion the difference between the rest energy and total relativistic energy was ignored due to its very small value in nonrelativistic interactions. The terms, $\beta_{j}$ and $\beta_{k}$, depend on the velocities of the two systems within the frame from which they are viewed. Relativistically, energy is transformed as the time component of the energy-momentum four-vector:
$E^{\prime}=\Gamma(E-up_{\parallel})$ (12)
where $E^{\prime}$ is the energy in the primed frame, $u$ is the velocity of the primed frame relative to the unprimed frame, $p_{\parallel}$ is the parallel component of momentum of the system in the unprimed frame, and capital $\Gamma$ is the relativistic transformation expression, $\frac{1}{\sqrt{1-\beta^{2}}}$ (where $\beta=u/c$). What must be shown is that the ratio of interaction energy, $\mathbf{V_{jk}}$, to the total relativistic energy (given in 11) is the same in all frames.
The transformation of interaction energy (represented as $\mathbf{V_{jk}}$) in a particular case depends on whether it contributes to the momentum in the unprimed frame. For example, in equation 12 above if $\mathbf{V_{jk}}$ is not a factor in the expression for $p_{\parallel}$ then the transformation is given simply by multiplication by $\Gamma$. This follows immediately from 11) and 12.
What we want to show is that, for an interaction between systems, $j$ and $k$, the ratio of the interaction energy to the total relativistic energy of the systems is the same, no matter what inertial frame the interaction is viewed from. In the rest frame of the center of mass of the interacting systems the total energy is given by equation 11 above. Since $\mathbf{p}_{j}=-\mathbf{p}_{k}$ in the c-o-m frame the total momentum is zero. Therefore, the transformation of both the interaction energy and the total relativistic energy from the c-o-m frame to any other frame is given simply by multiplication by $\Gamma$. Thus, the ratio is the same in every frame.
As described earlier the rate parameter associated with the stochastic operator, $\gamma_{jk}$, is designed to integrate to a value of order, $1$, over the period during which correlations are being established through the exchange of conserved quantities:
$\int\gamma_{jk}dt\;\sim\;1.$ (13)
[…]
$\int_{t}^{t+\Delta\,t}\,d\xi(s)d\xi^{*}(s)ds\;=\;\Delta\,t.$ (14)
Over the period of the interaction 13 and 14 imply that
$\int\sqrt{\gamma_{jk}}d\xi\;\sim\;1.$ (15)
It is also clear that transformations of the rate parameter, $\gamma_{jk}$, from one frame to another must vary in a synchronized manner relative to ${\mathcal{V}}_{jk}$. Therefore, the two key elements of the collapse term, ${\mathcal{V}}_{jk}$ and $\gamma_{jk}$, have the correct Lorentz transformation properties. [^8] It is true that the rate parameter, $\gamma_{jk}$, has been defined using the divergence of the current density based on the nonrelativistic Hamiltonian. However, for the low interaction energies and relative kinetic energies this yields a value essentially equivalent to the relativistic formula. Moreover, it varies inversely with respect to time according to the usual relativistic transformation formula. So, if we look at the primary stochastic term,
${\mathcal{V}}_{jk}\,\,\psi\,\sqrt{\gamma_{jk}}d\xi,$ (16)
everything transforms properly with the exception of the nonrelativistic version of the wave function, $\psi$. To complete the argument concerning Lorentz invariance we need to look at the entanglement relations that underlie the nonrelativistic approximation.
The fundamental idea behind this proposal is that wave functions are induced to collapse to one of their branches by the entangling interactions that constitute measurements. It is entanglement relations that define the branches and mediate collapse; the entanglement established by previous interactions is crucial to the collapse process.
This proposal was developed in a nonrelativistic framework in which the $j$ and $k$ systems are pictured as interacting instantaneously at spacelike separations. In reality the causal influences that generate the interactions propagate at finite speeds over timelike or lightlike intervals. In the simplified account based on the nonrelativistic framework both ends of an instantaneous jkj-k interaction are multiplied by the same ${\mathcal{V}}_{jk}\sqrt{\gamma_{jk}}d\xi$ term, and this is critical both in maintaining conservation laws and insuring the overall consistency of the collapse process. This is no longer true in the more realistic case involving finite propagation speeds; so we need to see how entanglement operates to maintain consistency.
Consider an interaction between systems $j$ and $k$. It consists of a sequence of causal influences emitted by one of the systems and the reaction of the other system to those influences. Suppose at time, $t_{1}$, system $j$ reacts to an influence emitted at a slightly earlier time by system $k$ and emits a causal influence. This interaction is accompanied by a stochastic change in the wave function proportional to ${\mathcal{V}}_{jk}\sqrt{\gamma_{jk}}d\xi\,(t_{1})$. The causal influence carries with it the effects of the stochastic operator at that time. At a slightly later time, $t_{2}$, system $k$ reacts to the influence from system $j$ and emits a causal influence. This interaction is also accompanied by a stochastic change proportional to ${\mathcal{V}}_{jk}\sqrt{\gamma_{jk}}d\xi\,(t_{2})$. Because of the entanglement that is generated by the interaction the effects of the stochastic operator acting at time, $t_{2}$, propagate back to system $j$. So both ends of the interaction are subject to the same two stochastic changes. The interaction proceeds in this “ping-pong” fashion. This ensures the consistency of the collapse process. The simplified, instantaneous interaction description gives a condensed, approximate version of this process.
This analysis can now be used to generate a quantitative estimate of the accuracy of the nonrelativistic approximation involving $\psi$. Because this proposal was formulated in a nonrelativistic framework the interaction energies are assumed to be small relative to the total relativistic energy of the systems involved. These interaction energies are typical of those used in a very wide range of measurements on quantum systems. Although the jkj-k system can have a very high energy in the preferred frame, the low interaction energies imply that in the center-of-mass frame the relevant velocities are nonrelativistic. Over the fairly small spacetime region in which the interaction occurs one can transform to the c-o-m frame and analyze the interaction with nonrelativistic quantum mechanics.
There are two primary factors that can affect the accuracy of the nonrelativistic approximation. The first of these stems from the fact that the nonrelativistic wave function, $\psi$, is defined on a different family of spacelike hypersurfaces from that of the preferred frame. The second concerns the extent to which the jkj-k components of the wave function, $\psi$, can evolve during the propagation delays. To evaluate the magnitude of these effects note first that the large majority of the amplitude transfer occurs close to the point at which the interaction energy reaches its maximum. This can be easily verified from the definitions of ${\mathcal{V}}_{jk}$ and $\gamma_{jk}$. Typical nonrelativistic measurement interactions take place over small distances and short times; fairly typical interactions can occur at $10^{-10}$ meters and $10^{-16}$ seconds. So we need to assess how large a fractional deviation from the correct amplitudes could be induced over this range.
To determine the maximum possible temporal discrepancy between various portions of the interacting jkj-k wave function associated with the hyperplane variation we can divide the interaction distance by the speed of light: $\frac{10^{-10}}{3*10^{8}}\;\sim\;10^{-18}$ seconds. Since the causal influences discussed above generally propagate at the speed of light discrepancies resulting from propagation delays are of the same magnitude. The maximum speeds for systems interacting at this (minimum) distance are about $10^{6}$ meters per second. So the maximum possible discrepancies in the positions of the various interacting portions of the systems are about $10^{-12}$ meters. This is about $0.01$ of the separation of the systems. This fraction provides a measure of accuracy for the approximation.
The considerations above apply to any interaction independent of the relationship of the preferred frame to the center-of-mass frame for that particular interaction. We have seen that the term, $\sqrt{\gamma_{jk}}d\xi$, integrates to a value of order $1$ no matter which frame is the preferred one. It is clear, therefore, that, independent of which reference frame is preferred, the average magnitude of the amplitude shift will be of order,
The direction and specific magnitude of the amplitude transfers vary from case to case, but this is determined by the random nature of the stochastic process - not by the selection of a preferred frame. Thus the amplitude shifts governed by 10 are Lorentz invariant.